an immersion heater has rating of 2kw,220v while in use. calculate
1) current passing through it
2)its resistance
1
Expert's answer
2012-06-05T10:30:27-0400
P = I · U = U^2 / R
As P = 220 kw, U = 220 v:
R = U^2 / P = 220 * 220 /
220000 = 0.22 Ohm.
I = P / U = 220000 / 220 = 1000 A.

Answer: 1 kA,
0.22 Ohm.